The comparative financial statements prepared at December 31 for Golden Corporation showed the following summarized data: Cur1-b. Are the current year results better, or worse, than those for the previous year? Better O Worse 2-a. Compute the net pro4-a. Stockholders equity totaled $42,000 at the beginning of the previous year. Compute the return on equity (ROE) ratios fo6-b. Is debt providing financing for a larger or smaller proportion of the companys asset growth? Larger Proportion Smaller
Answer #1

1-a) Gross profit percentage = Gross profit/sales x 100

Current year = 130,000/280,000 x 100 = 46.4%

Previous year = 105,000/245,000 x 100 = 42.9%

1-b) Better

2-a) Net profit margin = Net profit/sales x 100

Current year = 36,400/280,000 x 100 = 13%

Previous year = 29,000/245,000 x 100 = 11.8%

2-b) Better

3-a) Earnings per share = net income/number of common stock

No. of common stock = 42,000/5 = 8,400 shares

Current year = 36,400/8,400 = 4.33

Previous year = 29,000/8,400 = 3.45

3-b) Better

4-a) Return on equity = Net income/average shareholders equity x 100

Average shareholders equity of current year :-

= [(42,000 + 7,000 + 9,000) + (42,000 + 9,000 + 40,400)]/2 = 74,700

Average shareholders equity of previous year :-

= [(42,000) + (42,000 + 7,000 + 9,000)]/2 = 50,000

Current year = 36,400/74,700 x 100 = 48.7%

Previous year = 29,000/50,000 x 100 = 58%

4-b) Worse

5-a) Fixed assets turnover = Net sales/average fixed assets

Average fixed assets of current year = (65,000 + 58,000)/2 = 61,500

Average fixed assets of previous year = (58,000 + 45,000)/2 = 51,500

Current year = 280,000/61,500 = 4.55

Previous year = 245,000/51,500 = 4.76

5-b) Worse

6-a) Debt to assets ratio = total debt/total assets

Current year = (15,700 + 65,000)/172,100 = 0.47

Previous year = (33,000 + 65,000)/156,000 = 0.63

6-b) Smaller proportion

7-a) Times Interest earned = EBIT/Total interest expense

Current year EBIT = 130,000 - 73,300 = 56,700

Previous year EBIT = 105,000 - 66,400 = 38,600

Current year = 56,700/4,700 = 12.1

Previous year = 38,600/4,600 = 8.4

7-b) Better

8-a) price earnings ratio = MPS/EPS

Current year = 50/4.33 = 11.55

Previous year = 38/3.45 = 11.01

8-b) More optimistic